# Martian World Models: Controllable Video Synthesis with Physically Accurate 3D Reconstructions

### v1

@@ 1 Introduction
The advancement of space exploration is critically dependent on the development of robust robotic systems and operational procedures [12] tailored to diverse extraterrestrial environments. A major challenge across such domains is the lack of platforms capable of synthesizing realistic and dynamic data. This limitation hinders autonomous mission planning [37], operational rehearsal [61], rover navigation, and the execution of complex robotic tasks [21, 36]. Current publicly available extraterrestrial imagery, for example from NASA’s Martian rovers such as Curiosity and Perseverance—is typically provided as static stereo pairs captured from discrete and sparsely distributed viewpoints. The quality of these images is also affected by the interplanetary bandwidth limitations [14] and operational constraints. As a result, reconstruction of photorealistic 3D environments from such sparse and constrained imagery, a common challenge in planetary datasets, remains a obstacle.
Recent advances in video synthesis [6, 28, 24] offer promising avenues to mitigate these challenges. However, training an effective video generation model conditioned on sparse Martian imagery or reconstructed 3D models remains difficult. Most existing models [65, 55], which are typically trained on large-scale terrestrial datasets, struggle to generalize to the Martian domain due to the substantial domain gap. In addition, the challenges inherent to 3D reconstruction on Martian data often result in geometric models with insufficient accuracy, making it difficult to support high-fidelity, 3D-consistent video synthesis from such limited inputs.
@@ 2 Related works
Planetary Environment Simulation.   Prior efforts in simulating planetary environments [22, 53] have focused on enhancing the interpretation and interaction with Martian data. Approaches include Mixed Reality (MR) technologies [35, 38], such as those developed by NASA JPL’s Operations Lab and Microsoft [1, 4], enabling immersive exploration of 3D terrain models from rover data. Additionally, 3D reconstruction techniques like the MaRF [13] framework, which employs Neural Radiance Fields (NeRF) [39] for continuous volumetric representations from sparse images, have improved visualization from novel viewpoints. However, the MaRF framework is notably limited, having been demonstrated in different 3D environments. More broadly, these existing technologies often require high-quality, consistent visual data and exhibit limitations in scalability and adaptability. Our approach addresses these challenges by employing video generation models trained on a dedicated video dataset produced by our specialized data processing pipeline.
@@
3D Modeling from Sparse Views.   Reconstructing 3D scenes from sparse views [8] is a significant challenge. NeRF and 3DGS [26] typically demand hundreds of images and rely on the Structure-from-Motion (SfM) [46] approach (e.g., COLMAP [47]). To address this, some works leverage priors by pre-training on large datasets [7, 25, 66, 9, 23] or by applying regularization during NeRF optimization [56, 45, 49, 52, 51, 62, 34]. To mitigate the overfitting to input views in 3DGS, FSGS [71] and SparseGS [63] incorporate external priors from depth estimator with the optimization process. Others, like InstantSplat [11], utilize powerful 3D reconstruction models [30] to acquire accurate camera poses and initial geometries. However, robust sparse-view reconstruction remains an open problem, particularly in challenging environments such as the Martian surface, which exhibit textureless areas, repetitive patterns, and photometric variations. Our method addresses this gap by integrating a 3D geometric foundation model for initial geometric estimation with a specialized pipeline for refinement and neural scene representation from sparse stereo imagery, facilitating high-quality multimodal data synthesis.
@@
Conditional Generative Models.   Recent text-to-video models [65, 28, 55, 5] based on Diffusion Transformers (DiTs) [41] leverage the scalability of transformers, typically employing text encoders [42, 43], a 3D-VAE [67, 65] for video compression and tokenization, and a transformer generator that processes flattened video and text tokens. These architectures model spatiotemporal and textual information through global attention mechanisms or separate self- and cross-attention, leading to significant improvements in generation duration and temporal consistency. View-controllable video generation [60, 17, 33, 2, 68], which is crucial for immersive simulations, has seen efforts to integrate camera control into pretrained models. However, these methods, predominantly trained on standard terrestrial datasets, often exhibit difficulties with 3D consistency when applied to out-of-domain environments like Mars. In contrast, our work utilizes the M3arsSynth dataset, specifically curated with rich 3D geometric information, to train MarsGen, enabling physically plausible and precisely controllable Martian simulations with enhanced 3D consistency.
@@ 3.1 Martian Stereo Image Acquisition and Preprocessing
Systematic Data Filtering Strategy.   A multi-stage filtering pipeline ensures visual dataset integrity by systematically addressing imperfections. Firstly, low-resolution and grayscale images are eliminated. Thumbnails are discarded based on size heuristics. Grayscale images are removed based on low RGB channel variance. Secondly, redundant content is excluded using perceptual hashing [69]. This technique employs hash codes and Hamming distances to identify and remove near-duplicates captured under varied imaging conditions, preserving semantic diversity. Thirdly, blurry and low-sharpness images are rejected. A sharpness filter using Laplacian variance [3] discards images with low edge contrast to maintain geometric and photometric quality, which is vital for tasks like stereo reconstruction. Finally, frames exhibiting anomalous color distributions are filtered out. Irrelevant frames (e.g., obscured, malfunctions), identified by skewed or flat color intensity histograms, are removed to ensure clear, terrain-focused scenes for robust surface representation learning.(See Appendix A.1 for details.)
Semi-Automated Refinement.   To address complex visual artifacts like hardware occlusions and unfavorable lighting that degrade reconstruction quality (see Fig. 2), we employ a semi-automated refinement process. This step supplements manual image verification with Grounded-SAM [44], which we use to generate segmentation masks identifying non-terrain objects based on textual prompts. These masks guide human annotators to efficiently identify and discard compromised image data. The result is a curated collection of images with high visual integrity, providing a reliable foundation for robust stereo reconstruction. Further details on this preprocessing are provided in Appendix A.2.
@@ 3.2 Neural 3D Reconstruction from Navigation Stereo Cameras
Figure 3: Overview of the M3arsSynth dataset construction and conditional video generation through MarsGen. The red box outlines the data curation pipeline, the green box shows the obtained M3arsSynth dataset, and the blue box details our MarsGen model. We process stereo image pairs using a metric-aware foundation model and solve the Perspective-n-Point (PnP) [29] problem to reconstruct metric-scale 3D Martian scenes. Subsequently, video frames rendered from these scenes, together with text prompts and encoded camera trajectories, are then used to condition a Video Diffusion Transformer, enabling the synthesis of novel and controllable Martian video sequences.
@@
Camera Calibration.   Our 3D reconstruction pipeline operates on the standard pinhole camera model [15] and therefore requires the corresponding camera parameters. However, the PDS metadata accompanying our dataset lacks these crucial calibration parameters. To address this limitation, we infer the necessary parameters directly from the images. We employ the Visual Geometry Grounded Transformer (VGGT) [58], a feed-forward neural network designed to infer key 3D scene attributes from input views. Our empirical evaluations confirm that VGGT reliably predicts the camera intrinsics from the input stereo images, providing the essential parameters for subsequent reconstruction tasks.
@@
Dense 3D Geometry Initialization.   After obtaining camera intrinsics, we estimate metric deoth maps using a pretrained monocular depth estimation network [19] to construct dense 3D geometry. Each depth map is then back-projected using the inverse intrinsic matrix $K^{-1}$ to transform each pixel $(u,v)$ with depth $d$ into a 3D point in the corresponding camera coordinate system: $P_{c}=d\cdot K^{-1}[u,v,1]^{T}.$ This process yields initial per-view point clouds where each pixel has a direct 3D correspondence within its respective camera frame, providing a dense geometry initialization.
@@
Relative Pose Estimation and Geometric Refinement. To establish a coherent 3D model, we recover the relative camera extrinsics between the stereo images. We formulate this task as a Perspective-n-Point (PnP) [32] optimization, as PnP computes the camera pose from a set of 3D points and their corresponding 2D image projections. This approach utilizes robust 2D feature correspondences $\mathbf{p}_{1}\leftrightarrow\mathbf{p}_{2}$ between the stereo image pair, which are detected using the Generalizable Image Matcher (GIM) [50]. GIM’s notable generalization capability stems from its training on extensive and varied internet video data (approximately 180,000 image pairs from 50 hours of video) and its sophisticated self-training framework. This enables robust matching across diverse conditions and often surpasses traditional methods that rely on less diverse datasets or failure-prone 3D reconstruction processes [50]. Here, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ represent matched pixel coordinate vectors in the left and right views, respectively. Initial depth estimates from the monocular network are used to back-project points $\mathbf{p}_{1}$ into 3D space, yielding a set of 3D points $\mathbf{P}_{1}$. Given the known camera intrinsics $K$, these 2D-3D correspondences (formed by $\mathbf{P}_{1,i}$ and their corresponding $\mathbf{p}_{2,i}$) facilitate the estimation of the relative camera pose.
@@
The initial geometry derived from the monocular depth estimation may exhibit inconsistencies, particularly in scale across views. To ensure a coherent 3D reconstruction, we apply a depth rescaling step after pose estimation. Specifically, consider a set of $N$ corresponding 2D points $(\mathbf{p}_{1,j},\mathbf{p}_{2,j})$ between the two views, obtained via the GIM feature matching described above. Let $d_{1,j}$ and $d_{2,j}$ be their respective initial depth values. These depths are independently obtained by applying the monocular depth model to view 1 (for $\mathbf{p}_{1,j}$) and view 2 (for $\mathbf{p}_{2,j}$). We then seek to find a single scale factor $s$ and a bias $b$ that best align these corresponding depth measurements by solving the following least-squares regression problem:
$\min_{s,b}\sum_{j}(s\cdot d_{1,j}+b-d_{2,j})^{2}$ (2)
@@
3D Gaussian Splatting for Photorealistic Scene Modeling. Using the camera parameters and dense point cloud from prior stages, we optimize a 3DGS [26, 20] representation, initializing Gaussian primitives directly from the per-pixel points [11]. The optimization combines a photometric loss with a depth regularization loss [64, 63, 31, 27] that leverages our derived per-pixel correspondences to enforce geometric consistency. To mitigate significant appearance variations between Martian stereo pairs, often caused by differing camera settings or lighting, we integrate a bilateral grid [59] into the process. Furthermore, since standard k-nearest neighbor initialization can produce overly large Gaussian scales and degrade depth accuracy, we adopt a modified strategy [10] where scale is determined by the nearest point along the depth axis.
@@ 3.3 Multimodal Martian Dataset Creation
Video, Normal, Trajectory, Depth, Text. From a set of adaptively scaled camera trajectories, we generate a comprehensive multimodal dataset. Novel view videos, accompanied by their corresponding depth and normal maps, are produced by rendering the scene along these trajectories. The Trajectory modality encompasses the precise 6-DOF pose parameters for each frame ($T_{t}$), from which textual descriptions of camera motion are subsequently derived. To capture scene content, we further generate textual captions by applying a Vision Language Model (VLM) [40] to the rendered videos. This procedure results in a rich dataset comprising visual, geometric, and textual modalities, all structured around the foundational camera trajectories.
@@ 3.4 Metrics Evaluating 3D Consistency
To assess the 3D consistency of video sequences generated by our MarsGen model (trained on the M3arsSynth dataset), we employ the 2D Warp Error metric [57].
@@ 4 Martian Terrain Video Generator Training
Model Architecture. MarsGen is built upon the Video Diffusion Transformer (VDiT) framework [41]. The video and text prompts are first encoded into a latent space [67, 43] and concatenated. A 3D self-attention mechanism, operating across both temporal and spatial dimensions, facilitates a comprehensive interaction between the multimodal information. Finally, a decoder reconstructs the latent representations back into the video space. We train this model on the M3arsSynth dataset to adapt it to the unique visual cues and structural characteristics of Martian environments, thus supporting high-fidelity and controllable video generation under domain-specific constraints.
Controllable Content Generation. To achieve fine-grained control over the generative process, MarsGen incorporates multiple modalities, including a reference initial frame, textual prompts (natural language descriptions), and camera trajectories. These inputs jointly guide the model across spatial, temporal and semantic dimensions. To circumvent the computational demands of full model fine-tuning, we employ a lightweight conditioning strategy inspired by ControlNet architecture [70]. Specifically, we inject control signals into intermediate layers of the pretrained VDiT backbone, enabling the modulation of generation dynamics without disrupting its learned visual priors.
Video Model Training. We initialize the VDiT backbone with pretrained weights from CogVideoX-5B-I2V [67, 65]. We then fine-tune the model, incorporating its controllable branch, on 8 A100 GPUs for 8,000 iterations using the M3arsSynth dataset.
@@
Pyramidal-Flow [24] 78.495 637.952 17.930 -- -- --
CogVideoX [65] 48.912 411.808 6.866 -- -- --
Kling 74.632 727.130 24.793 -- -- --
Sora [6] 142.954 823.418 10707.713 -- -- --
@@
CameraCtrl [16] 123.386 772.476 17.410 20.014 0.288 0.408
ViewCrafter [68] 169.942 2297.899 501.734 13.143 0.262 0.586
Ours 38.779 364.822 6.071 21.239 0.336 0.351
Table 2: Quantitative comparison of reconstruction pipelines for Martian stereo imagery. We compare COLMAP, the Transformer-based MASt3R [30], and our metric-aware initialization method across four key metrics: runtime efficiency (Time), frame-level robustness (Data Utilization), geometric accuracy (2D Reprojection Error [57]), and reconstruction density (Point Number). Our pipeline fully utilizes all available data and achieves competitive accuracy, whereas COLMAP fails on nearly 30% of the preprocessed image pairs.
@@
COLMAP [46, 48] 71.8 3.6 0.134 $\sim 2,000$
MASt3R [30] 100.0 8.7 46.98 $\sim 250,000$

### v2

# Martian World Model: Controllable Video Synthesis with Physically Accurate 3D Reconstructions

@@ 1 Introduction
The advancement of space exploration is critically dependent on the development of robust robotic systems and operational procedures gao2017review tailored to diverse extraterrestrial environments. A major challenge across such domains is the lack of platforms capable of synthesizing realistic and dynamic data. This limitation hinders autonomous mission planning maurette2003mars, operational rehearsal wright2005terrain, rover navigation, and the execution of complex robotic tasks huntsberger2000robotics; mathers2012robotic. Current publicly available extraterrestrial imagery, for example from NASA’s Martian rovers such as Curiosity and Perseverance—is typically provided as static stereo pairs captured from discrete and sparsely distributed viewpoints. The quality of these images is also affected by the interplanetary bandwidth limitations 5408994 and operational constraints. As a result, reconstruction of photorealistic 3D environments from such sparse and constrained imagery, a common challenge in planetary datasets, remains a obstacle.
Recent advances in video synthesis videoworldsimulators2024; kong2024hunyuanvideo; jin2024pyramidal offer promising avenues to mitigate these challenges. However, training an effective video generation model conditioned on sparse Martian imagery or reconstructed 3D models remains difficult. Most existing models yang2024cogvideox; wang2025wan, which are typically trained on large-scale terrestrial datasets, struggle to generalize to the Martian domain due to the substantial domain gap. In addition, the challenges inherent to 3D reconstruction on Martian data often result in geometric models with insufficient accuracy, making it difficult to support high-fidelity, 3D-consistent video synthesis from such limited inputs.
@@ 2 Related works
Planetary Environment Simulation.   Prior efforts in simulating planetary environments jain2003roams; tian20243d have focused on enhancing the interpretation and interaction with Martian data. Approaches include Mixed Reality (MR) technologies mahmood2019improving; memarsadeghi2020virtual, such as those developed by NASA JPL’s Operations Lab and Microsoft abercrombie2017onsight; beaton2020mission, enabling immersive exploration of 3D terrain models from rover data. Additionally, 3D reconstruction techniques like the MaRF giusti2022marf framework, which employs Neural Radiance Fields (NeRF) mildenhall2021nerf for continuous volumetric representations from sparse images, have improved visualization from novel viewpoints. However, the MaRF framework is notably limited, having been demonstrated in different 3D environments. More broadly, these existing technologies often require high-quality, consistent visual data and exhibit limitations in scalability and adaptability. Our approach addresses these challenges by employing video generation models trained on a dedicated video dataset produced by our specialized data processing pipeline.
@@
3D Modeling from Sparse Views.   Reconstructing 3D scenes from sparse views chen2024survey is a significant challenge. NeRF and 3DGS 3dgs typically demand hundreds of images and rely on the Structure-from-Motion (SfM) schoenberger2016sfm approach (e.g., COLMAP schonberger2016structure). To address this, some works leverage priors by pre-training on large datasets chen2021mvsnerf; johari2022geonerf; yu2021pixelnerf; chibane2021stereo; jang2021codenerf or by applying regularization during NeRF optimization wang2023sparsenerf; roessle2023ganerf; seo2023flipnerf; somraj2023simplenerf; somraj2023vip; wynn2023diffusionerf; liu2024deceptive. To mitigate the overfitting to input views in 3DGS, FSGS zhu2024fsgs and SparseGS xiong2024sparsegs incorporate external priors from depth estimator with the optimization process. Others, like InstantSplat fan2024instantsplat, utilize powerful 3D reconstruction models mast3r to acquire accurate camera poses and initial geometries. However, robust sparse-view reconstruction remains an open problem, particularly in challenging environments such as the Martian surface, which exhibit textureless areas, repetitive patterns, and photometric variations. Our method addresses this gap by integrating a 3D geometric foundation model for initial geometric estimation with a specialized pipeline for refinement and neural scene representation from sparse stereo imagery, facilitating high-quality multimodal data synthesis.
@@
Conditional Generative Models.   Recent text-to-video models yang2024cogvideox; kong2024hunyuanvideo; wang2025wan; blattmann2023stable based on Diffusion Transformers (DiTs) peebles2023scalable leverage the scalability of transformers, typically employing text encoders radford2021learning; raffel2020exploring, a 3D-VAE yu2023language; yang2024cogvideox for video compression and tokenization, and a transformer generator that processes flattened video and text tokens. These architectures model spatiotemporal and textual information through global attention mechanisms or separate self- and cross-attention, leading to significant improvements in generation duration and temporal consistency. View-controllable video generation wang2024motionctrl; he2024cameractrl; liang2024wonderland; bahmani2024ac3d; yu2024viewcrafter, which is crucial for immersive simulations, has seen efforts to integrate camera control into pretrained models. However, these methods, predominantly trained on standard terrestrial datasets, often exhibit difficulties with 3D consistency when applied to out-of-domain environments like Mars. In contrast, our work utilizes the M3arsSynth dataset, specifically curated with rich 3D geometric information, to train MarsGen, enabling physically plausible and precisely controllable Martian simulations with enhanced 3D consistency.
@@ 3.1 Martian Stereo Image Acquisition and Preprocessing
Systematic Data Filtering Strategy.   A multi-stage filtering pipeline ensures visual dataset integrity by systematically addressing imperfections. Firstly, low-resolution and grayscale images are eliminated. Thumbnails are discarded based on size heuristics. Grayscale images are removed based on low RGB channel variance. Secondly, redundant content is excluded using perceptual hashing zauner2010implementation. This technique employs hash codes and Hamming distances to identify and remove near-duplicates captured under varied imaging conditions, preserving semantic diversity. Thirdly, blurry and low-sharpness images are rejected. A sharpness filter using Laplacian variance bansal2016blur discards images with low edge contrast to maintain geometric and photometric quality, which is vital for tasks like stereo reconstruction. Finally, frames exhibiting anomalous color distributions are filtered out. Irrelevant frames (e.g., obscured, malfunctions), identified by skewed or flat color intensity histograms, are removed to ensure clear, terrain-focused scenes for robust surface representation learning.(See Appendix A.1 for details.)
Semi-Automated Refinement.   To address complex visual artifacts like hardware occlusions and unfavorable lighting that degrade reconstruction quality (see Fig. 2), we employ a semi-automated refinement process. This step supplements manual image verification with Grounded-SAM ren2024grounded, which we use to generate segmentation masks identifying non-terrain objects based on textual prompts. These masks guide human annotators to efficiently identify and discard compromised image data. The result is a curated collection of images with high visual integrity, providing a reliable foundation for robust stereo reconstruction. Further details on this preprocessing are provided in Appendix A.2.
@@ 3.2 Neural 3D Reconstruction from Navigation Stereo Cameras
Figure 3: Overview of the M3arsSynth dataset construction and conditional video generation through MarsGen. The red box outlines the data curation pipeline, the green box shows the obtained M3arsSynth dataset, and the blue box details our MarsGen model. We process stereo image pairs using a metric-aware foundation model and solve the Perspective-n-Point (PnP) lepetit2009epnp problem to reconstruct metric-scale 3D Martian scenes. Subsequently, video frames rendered from these scenes, together with text prompts and encoded camera trajectories, are then used to condition a Video Diffusion Transformer, enabling the synthesis of novel and controllable Martian video sequences.
@@
Camera Calibration.   Our 3D reconstruction pipeline operates on the standard pinhole camera model hartley2003multiple and therefore requires the corresponding camera parameters. However, the PDS metadata accompanying our dataset lacks these crucial calibration parameters. To address this limitation, we infer the necessary parameters directly from the images. We employ the Visual Geometry Grounded Transformer (VGGT) wang2025vggt, a feed-forward neural network designed to infer key 3D scene attributes from input views. Our empirical evaluations confirm that VGGT reliably predicts the camera intrinsics from the input stereo images, providing the essential parameters for subsequent reconstruction tasks.
@@
Dense 3D Geometry Initialization.   After obtaining camera intrinsics, we estimate metric deoth maps using a pretrained monocular depth estimation network hu2024metric3d to construct dense 3D geometry. Each depth map is then back-projected using the inverse intrinsic matrix $K^{-1}$ to transform each pixel $(u,v)$ with depth $d$ into a 3D point in the corresponding camera coordinate system: $P_{c}=d\cdot K^{-1}[u,v,1]^{T}.$ This process yields initial per-view point clouds where each pixel has a direct 3D correspondence within its respective camera frame, providing a dense geometry initialization.
@@
Relative Pose Estimation and Geometric Refinement. To establish a coherent 3D model, we recover the relative camera extrinsics between the stereo images. We formulate this task as a Perspective-n-Point (PnP) li2012robust optimization, as PnP computes the camera pose from a set of 3D points and their corresponding 2D image projections. This approach utilizes robust 2D feature correspondences $\mathbf{p}_{1}\leftrightarrow\mathbf{p}_{2}$ between the stereo image pair, which are detected using the Generalizable Image Matcher (GIM) shen2024gim. GIM’s notable generalization capability stems from its training on extensive and varied internet video data (approximately 180,000 image pairs from 50 hours of video) and its sophisticated self-training framework. This enables robust matching across diverse conditions and often surpasses traditional methods that rely on less diverse datasets or failure-prone 3D reconstruction processes shen2024gim. Here, $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ represent matched pixel coordinate vectors in the left and right views, respectively. Initial depth estimates from the monocular network are used to back-project points $\mathbf{p}_{1}$ into 3D space, yielding a set of 3D points $\mathbf{P}_{1}$. Given the known camera intrinsics $K$, these 2D-3D correspondences (formed by $\mathbf{P}_{1,i}$ and their corresponding $\mathbf{p}_{2,i}$) facilitate the estimation of the relative camera pose.
@@
The initial geometry derived from the monocular depth estimation may exhibit inconsistencies, particularly in scale across views. To ensure a coherent 3D reconstruction, we apply a depth rescaling step after pose estimation. Specifically, we first back-project the depth map $D_{0}$ from view 0 to reconstruct its corresponding 3D point cloud $P_{0}$. Then we project it onto view 1’s image plane. This warping process yields a new sparse depth map $D^{\prime}_{0\to 1}$ and a Boolean mask $M_{sparse}$, indicating valid projection areas in view 1. The depth value $d^{\prime}_{1}$ is stored in $D^{\prime}_{0\to 1}$ at the projected pixel coordinates $(u^{\prime}_{1}/d^{\prime}_{1},v^{\prime}_{1}/d^{\prime}_{1})$.
$P_{0}=D_{0}(p_{0})\cdot K_{0}^{-1}\begin{bmatrix}u_{0}\\ v_{0}\\ 1\end{bmatrix}$
$\begin{bmatrix}u^{\prime}_{1}\\ v^{\prime}_{1}\\ d^{\prime}_{1}\end{bmatrix}=K_{1}(RP_{0}+t)$
Then, we align the original monocular depth map of view 1 $D_{1}$, with the warped sparse depth map $D^{\prime}_{0\to 1}$. We solve a least-squares regression problem only within the valid region defined by the mask $M_{sparse}$.
$\min_{s,b}\sum_{(u,v)\in M_{sparse}}(s\cdot D_{1}(u,v)+b-D^{\prime}_{0\to 1}(u,v))^{2}$
@@
3D Gaussian Splatting for Photorealistic Scene Modeling. Using the camera parameters and dense point cloud from prior stages, we optimize a 3DGS 3dgs; huang20242d representation, initializing Gaussian primitives directly from the per-pixel points fan2024instantsplat. The optimization combines a photometric loss with a depth regularization loss xu2025depthsplat; xiong2024sparsegs; li2024dngaussian; kerbl2024hierarchical that leverages our derived per-pixel correspondences to enforce geometric consistency. To mitigate significant appearance variations between Martian stereo pairs, often caused by differing camera settings or lighting, we integrate a bilateral grid wang2024bilateral into the process. Specifically, our implementation is based on gsplat ye2025gsplat. We apply a per-view 3D bilateral grid as a differentiable post-processing layer to the rendered image, which models view-dependent effects. This grid is jointly optimized with the Gaussian parameters by minimizing the difference between the post-processed render and the corresponding training view, while a total variation loss is used to regularize the grid for smoothness. Furthermore, since standard k-nearest neighbor initialization can produce overly large Gaussian scales and degrade depth accuracy, we adopt a modified strategy cong2025videolifter where scale is determined by the nearest point along the depth axis.
@@ 3.3 Multimodal Martian Dataset Creation
Video, Normal, Trajectory, Depth, Text. From a set of adaptively scaled camera trajectories, we generate a comprehensive multimodal dataset. Novel view videos, accompanied by their corresponding depth and normal maps, are produced by rendering the scene along these trajectories. The Trajectory modality encompasses the precise 6-DOF pose parameters for each frame ($T_{t}$), from which textual descriptions of camera motion are subsequently derived. To capture scene content, we further generate textual captions by applying a Vision Language Model (VLM) openai2024chatgpt4o to the rendered videos. This procedure results in a rich dataset comprising visual, geometric, and textual modalities, all structured around the foundational camera trajectories.
@@ 3.4 Metrics Evaluating 3D Consistency
To assess the 3D consistency of video sequences generated by our MarsGen model (trained on the M3arsSynth dataset), we employ the 2D Warp Error metric wang2024vggsfm.
@@ 4 Martian Terrain Video Generator Training
Model Architecture. MarsGen is built upon the Video Diffusion Transformer (VDiT) framework peebles2023scalable. The video and text prompts are first encoded into a latent space yu2023language; raffel2020exploring and concatenated. A 3D self-attention mechanism, operating across both temporal and spatial dimensions, facilitates a comprehensive interaction between the multimodal information. Finally, a decoder reconstructs the latent representations back into the video space. We train this model on the M3arsSynth dataset to adapt it to the unique visual cues and structural characteristics of Martian environments, thus supporting high-fidelity and controllable video generation under domain-specific constraints.
Controllable Content Generation. To achieve fine-grained control over the generative process, MarsGen incorporates multiple modalities, including a reference initial frame, textual prompts (natural language descriptions), and camera trajectories. These inputs jointly guide the model across spatial, temporal and semantic dimensions. To circumvent the computational demands of full model fine-tuning, we employ a lightweight conditioning strategy inspired by ControlNet architecture zhang2023adding. Specifically, we inject control signals into intermediate layers of the pretrained VDiT backbone, enabling the modulation of generation dynamics without disrupting its learned visual priors.
Video Model Training. We initialize the VDiT backbone with pretrained weights from CogVideoX-5B-I2V yu2023language; yang2024cogvideox. We then fine-tune the model, incorporating its controllable branch, on 8 A100 GPUs for 8,000 iterations using the M3arsSynth dataset.
@@
Pyramidal-Flow jin2024pyramidal 78.495 637.952 17.930 - - -
CogVideoX yang2024cogvideox 48.912 411.808 6.866 - - -
Kling 74.632 727.130 24.793 - - -
Sora videoworldsimulators2024 142.954 823.418 10707.713 - - -
@@
CameraCtrl hecameractrl 123.386 772.476 17.410 20.014 0.441 0.408
ViewCrafter yu2024viewcrafter 169.942 2297.899 501.734 13.143 0.500 0.586
Ours 38.779 364.822 6.071 21.239 0.614 0.351
Table 2: Quantitative comparison of reconstruction pipelines for Martian stereo imagery. We compare COLMAP, the Transformer-based MASt3R mast3r, and our metric-aware initialization method across four key metrics: runtime efficiency (Time), frame-level robustness (Data Utilization), geometric accuracy (2D Reprojection Error wang2024vggsfm), and reconstruction density (Point Number). Our pipeline fully utilizes all available data and achieves competitive accuracy, whereas COLMAP fails on nearly 30% of the preprocessed image pairs.
@@
COLMAP schoenberger2016sfm; schoenberger2016mvs 71.8 3.6 0.134 $\sim 2,000$
MASt3R mast3r 100.0 8.7 46.98 $\sim 250,000$